OpenStudy (anonymous):

I am currently taking a practice test. The question is to find the x and y intercepts of the graph y=3x^2+6x-9. They say that the answer is x=-3 or x=2 and y=(0,-9). I get that x=1 or x=-3 and y=-9. What did i do wrong?

sam (.sam.):

y-intercept when x=0 y=-9 (0,9)

OpenStudy (anonymous):

Right.

sam (.sam.):

x-intercept when y=0 x=-3,1 (-3,0) and (1,0)

OpenStudy (across):

You are given \(f(x)=3x^2+6x-9\). To find the \(x\)-intercepts, you have to set \(3x^2+6x-9=0\implies3(x-1)(x+3)=0\implies x=1\text{ and }x=-3\). To find the \(y\)-intercept, you evaluate the function at \(0\): \(f(0)=-9\).

OpenStudy (anonymous):

Thank You. I thought that I was going crazy... Thank you!